Opened 2 years ago

Closed 21 months ago

#620 closed flight processing (fixed)

ES17/136, flight day 167c/2017, Albacete

Reported by: asm Owned by:
Priority: immediate Milestone: The Glorious Future
Component: Processing: general Keywords:
Cc: Other processors:

Description

Data location: ~arsf/arsf_data/2017/flight_data/spain/ES17_136-2017_167c_Albacete

Data arrived from NERC-ARF via hard drive 2017-06-30

Scientific objective: Develop LLT/SSE using AisaOwl data
Priority: Urgent

PI: Martin Wooster

Sensors:

  • Fenix (flown, requested)
  • Leica LIDAR (flown, not requested)
  • OWL (flown, requested)
  • PhaseOne (flown, requested)

Change History (27)

comment:1 Changed 2 years ago by asm

Unpacking

Project has been unpacked. There are 6 fenix flightlines and 6 owl lines.

Checks have been run:
-> Less than usual timestamps for owl flightline 5 nav file
-> 2 dropped frames in the owl files
-> 2 lidar files are empty:170616_171556/LDR170616_171556_0000000008.scn and 170616_171132/LDR170616_171132_0000000008.scn

comment:2 Changed 2 years ago by dac

dGPS data collected at the same time as flight provided by PI and copied to: /users/rsg/arsf/arsf_data/2017/misc/basestation_data/ES17_136_Basestation_Data

comment:3 Changed 2 years ago by dac

Digital Camera

Converted raw images to tiff.

comment:4 Changed 2 years ago by lah

Owl

Processed to level 1b to compare temperature with 167a and 167b.

comment:5 Changed 2 years ago by lah

Navigation - Basestation

Lat: 39 03 37.37207
Lon: -2 05 53.40719
Height: 748.661 m

comment:6 Changed 2 years ago by lah

Navigation Processing

Best position separation obtain with default configuration settings. Cropping the data to just the flight lines gives a better position separation, but still a bit spikey in places..

comment:7 Changed 2 years ago by lah

Owl Processing

All lines look good with sct = 2.00. Proceeding to map final files.

comment:8 Changed 2 years ago by lah

Owl Processing

Delivery made - Ready for delivery check.

comment:9 Changed 2 years ago by dac

Hyperspectral processing

Set three band files running to check SCT values.

comment:10 Changed 2 years ago by dac

Hyperspectral processing

Found SCT values:

Flightline FENIX
1 2
2 2
3 1
4 1
5 2
6 2

comment:11 Changed 2 years ago by dac

Basestation

NERC GEF retrieved basestation height using Leica GeoOffice. Total height 1.328 m. Will reprocess navigation data with correct height.

comment:12 Changed 2 years ago by dac

Fenix

Reprocessed navigation data with correct basestation height.

Ran through Fenix data again, using view vectors for loan sensor.

comment:13 Changed 2 years ago by dac

Digital Camera Processing

First 16 images didn't match with GPS. Removed from events file.

comment:14 Changed 2 years ago by dac

Digital Camera Processing

Files tagged and thumbnails generated.

comment:15 Changed 2 years ago by lah

Owl reprocessing

Looks like Owl was sent to reprocess without the level 1b files, so files were not mapped. Have sent level 1b files to reprocess as it's easier than renaming the delivery files.

comment:16 Changed 2 years ago by lah

Owl ready for Delivery Check

Fixed problems with scripts related to the mask and finished creating the delivery. Ready for checking.

comment:17 Changed 2 years ago by dac

Owl delivery check

  • Readme still contains picture of Dornier - removed
  • Removed LiDAR from logsheet and tidied up a little
  • check_apl_cmd ran through OK

Marking as ready to deliver and zipping mapped files

comment:18 Changed 2 years ago by dac

Hyperspectral processing

Ready for delivery check.

comment:19 Changed 2 years ago by lah

Fenix Delivery Check

Everything looks ok so marking as ready to deliver and zipping mapped files.

comment:20 Changed 2 years ago by dac

Data delivery

Fenix and Owl data placed on FTP server (slot 4). Notification sent to PI.

comment:21 Changed 22 months ago by asm

Digital Camera

Created delivery and readme file. Project ready for delivery check

comment:22 Changed 22 months ago by dac

Digital Camera Delivery Check

Removed LiDAR from KML. Other than that looks good. Ready to deliver.

comment:23 Changed 22 months ago by asm

Camera Delivery

Placed on FTP server (slot 4). Notification sent to PI on 18/10/2017.

comment:24 Changed 21 months ago by asm

Archiving

Tidying up before archive. Will delete Owl temperature directory as the owl sct problem was fixed and files can be recreated.

comment:25 Changed 21 months ago by asm

Archiving

Started uploading data to CEDA.

comment:26 Changed 21 months ago by asm

Archiving

Paused the upload of data to CEDA.

comment:27 Changed 21 months ago by dac

  • Resolution set to fixed
  • Status changed from new to closed
Note: See TracTickets for help on using tickets.